What is the potential energy of two charges of +4.2 uC and +6.1 uC which are separated by the distance of 50.0 cm

1
Expert's answer
2022-03-03T12:05:16-0500

The potential energies between two charges is

"U = \dfrac{kq_1q_2}{r} = \dfrac{9\cdot10^9\,\mathrm{N\cdot m^2/C^2}\cdot4.2\cdot10^{-6}\,\mathrm{C}\cdot6.1\cdot10^{-6}\,\mathrm{C}}{0.5\,\mathrm{m}} , \\
U = 0.46\,\mathrm{J}."

The energy is positive: two positive charges repel from each other.